Five gallons of a dark green latex paint that is 20% yellow paint are combined with a lighter green latex paint that is 40% yellow paint. How many galloons of the lighter green paint must be used to create a green paint that is 25% yellow paint?
.20x+.40(x+5)=1.25
.20x+.40x+2=1.25
.60x+2=1.25
.60x=.75
x=1.083????

Answer by Alan3354(69443)

Five gallons of a dark green latex paint that is 20% yellow paint are combined with a lighter green latex paint that is 40% yellow paint. How many galloons of the lighter green paint must be used to create a green paint that is 25% yellow paint?
.20x+.40(x+5)=1.25
.20x+.40x+2=1.25
.60x+2=1.25
.60x=.75
x=1.083????
=====================
5*20 + x*40 = (5+x)*25